OpenStudy (anonymous):

(h o h)(-4); Let g(x)=2x and h(x)=x^2+4. Evaluate each expression. How do I?

OpenStudy (anonymous):

is it hoh, or hog, goh?

OpenStudy (anonymous):

Its hoh

OpenStudy (anonymous):

Okay, you just have to compose the function with itself. With composition, you know that\[fog(x)=f(g(x))\]so for hoh, it's just\[hoh(x)=h(h(x))=(h(x))^4+4=(x^2+4)^4+4\]

OpenStudy (anonymous):

You can then just find (h o h)(-4)

OpenStudy (anonymous):

(16+4)^4+4 = 20^4+4=160,004

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

become a fan, jaiquan ;)

OpenStudy (anonymous):

how do I become a fan? New to this.

OpenStudy (anonymous):

There should be a blue link next to my name saying, "Become a fan". If it doesn't, refresh the page.